OpenStudy (anonymous):

solve and show work! (1/6b^2) + (1/6b) = (1/b^2)

OpenStudy (anonymous):

(1/6b^2) + (1/6b) = (1/b^2) multiply both sides of the equation by 1/b to get (1/6b) + (1/6) = (1/b) then subtract both sides by (1/b) to get (1/6b) + (1/6)-(1/b) = 0 then subtract both sides by (1/6) to get (1/6b) - (1/b) = -(1/6) now solve the simplify the left side to one fraction (basically subtract the fractions) to get (-5b/6b^2) = -(1/6) this simplifies to (-5/6b) = -(1/6) multiply both sides by 6b to get -5 = -b so 5 = b is your solution plug it in and you get 1/(6*25) + 1/(30) = 1/ 25 ----> 1/25 = 1/25